### v1

# Quantum Field Theory on Multifractal Spacetime: Varying Dimension and Ultraviolet Completeness

Quantum Field Theory on Multifractal Spacetime: Varying Dimension and Ultraviolet Completeness
Alessio Maiezza1,2∗, Juan Carlos Vasquez3†
1Dipartimento di Scienze Fisiche e Chimiche, Università
degli Studi dell’Aquila, via Vetoio, I-67100, L’Aquila, Italy,
2INFN, Laboratori Nazionali del Gran Sasso, 67010 Assergi, L’Aquila, Italy,
3Centro de Fisica Fundamental, Universidad de Los Andes, Merida, Venezuela
E-mail: alessiomaiezza@gmail.com, juancarlos8866@gmail.com
Inspired by various quantum gravity approaches, we explore quantum field theory where spacetime exhibits scaling properties and dimensional reduction with changing energy scales, effectively behaving as a multifractal manifold. Working within canonical quantization, we demonstrate how to properly quantize fields in such multifractal spacetime. Our analysis reveals that a non-differentiable nature of spacetime is not merely compatible with quantum field theory but significantly enhances its mathematical foundation. Most notably, this approach guarantees the finiteness of the theory at all orders in perturbation theory and enables rigorous construction of the $S$-matrix in the interaction picture. The multifractal structure tames dominant, large-order divergence sources in the perturbative series and resolves the Landau pole problem through asymptotic safety, substantially improving the theory’s behavior in the deep ultraviolet regime. Our formulation preserves all established predictions of standard quantum field theory at low energies while offering novel physical behaviors at high energy scales.
1 Introduction
While the concept of dimension spans broadly across geometry, physics often simplifies it by stating that we inhabit a three-dimensional space. Our macroscopic experience confirms this, yet exceptions may emerge at the microscopic level. Mathematics offers various definitions of dimension – the topological dimension $d$ represents an integer, with $d=3$ identified as the dimension of our physical space. In contrast, certain definitions do not require integer values, such as the Hausdorff dimension, which becomes essential for describing fractal manifolds [1].
Consider a sheet of paper, modeled topologically as two-dimensional. When crumpled, however, it assumes a different effective fractal dimension. We can estimate this physically by measuring the crumpled paper’s density. Solid objects have densities scaling as $r^{-3}$ (where $r$ represents a characteristic length), yielding $d=3$ as a dimensionality estimate that coincides with the topological dimension. The crumpled paper, however, follows a different scaling relation, typically $r^{-d^{eff}}$, where $32<d^{eff}<$ 3. This simple example illustrates a crucial insight: the physical dimension may be neither trivial nor externally fixed. Rather, it requires a “physical estimator”, a concept central to this work, analogous to the density in our example.
Such estimators become vital in quantum realms, where direct dimensional experience is impossible. Quantum field theory (QFT) provides a natural dimension estimator through the Green function. In Euclidean space, this function diverges as $r^{-(d-2)}$ at short distances. Conventionally, dimension is treated as a fixed external parameter conditioned by macroscopic experience. However, we can invert this logic: we examine Green function behavior at short distances to estimate the spacetime dimension  [^1] The concept that the dimension must be a posteriori determined was introduced in [2], specifically via a variational principle.. This transforms dimension into an observable quantity that may differ from the topological dimension $d$, an approach we adopt throughout this paper.
How should dimension differ from large-scale observations? Space-time could possess intrinsic fractal properties, causing fields to experience a different dimension than the topological one. The idea of a rough spacetime was first proposed by Wheeler who called it spacetime foam [3] – see also [4, 5]. We shall broadly refer to the roughness (non-differentiability) of spacetime as “fractal spacetime”. While string theory often suggests topological dimensions exceeding four (albeit warped), many quantum gravity approaches propose the opposite: dimensional reduction. Theories including Causal Dynamical Triangulations [6, 7, 8], Asymptotic Safety [9, 10], Causal Set Theory [11], and Loop Quantum Gravity [12] suggest that effective dimension flows continuously from four at large scales to two at short distances. Additional perspectives appear in [13, 14]. A model universe with variable dimension is presented in [15].
This article examines this phenomenon from the QFT perspective, broadening the seminal work of [16] on QFT formulation in fractal spacetime in many directions. The propagator, central to QFT, can serve as a dimensional estimator through running couplings. We investigate how canonical quantization should be modified to accommodate fractal space-time and dimensional reduction. Our findings reveal a straightforward generalization of canonical quantization that significantly enhances QFT consistency. First, the theory becomes finite at all orders in perturbation theory. Second, the behavior of the perturbative series is improved since the leading singularities in the Borel plane, namely, the ultraviolet (UV) renormalons, disappear. Third, also the UV Landau pole disappears so that any QFT model embedded in multifractal spacetime becomes asymptotically safe. We trace these improvements back to the Poincaré non-invariance of the vacuum, enabling robust $S$-matrix construction by circumventing limitations imposed by the Haag theorem [17]. The resulting theory agrees with the standard quantum field theory at low energies, while all the changes and improvements would appear at some high energy scale we denote $M$.
In Sec. 2, we provide a self-contained analysis of how multifractality may be connected to dimensional reduction, consistent with the quantum gravity approaches mentioned above. We address canonical quantization in Sec. 3, then we discuss the consistency in Sec. 4, and the implications in Sec. 5. We give our outlook in Sec. 6. Additional technical details are reported in the appendix A.
2 Fractality, smoothing and dimensional reduction
This preliminary section aims to connect the concepts of multifractal space and varying effective dimensions.
To start with, consider first a 1D space where the tangent is not defined at any point, namely, is not smooth, and consider a non-differentiable function $f(x):\mathcal{M}\rightarrow\mathbb{R}$ on this 1D manifold $\mathcal{M}$. It is possible to build from $f$ a new function $F:\mathcal{M}\rightarrow\mathbb{R}$, differentiable, via convolution with an auxiliary smoothing function $h(x)$ that can be chosen Gaussian[18],
$h(x)=\frac{2}{\sqrt{4\pi l^{2}}}e^{-\frac{x^{2}}{4l^{2}}}\,.$ (1)
The convolution product $f*h$ gives
[…]
now being $F$ smoothed, differentiable, and dependent on an intrinsic scale $l$, implying resolution dependence. By construction, when $l\rightarrow 0$, the function $F$ coincides with the original non-differentiable function $f(x)$.
For instance, think of the above function $f(x)$ as the distance between a horizontal line and a point in the 1D non-differentiable line given by the points belonging to the Koch curve. The function $f$ is continuous everywhere in $x$ but non-differentiable anywhere in $x$ [19]. Now assume that one is interested in finding the distance between any two points in the Koch curve. For zero resolution, such distance is not well-defined (infinite) when $l\rightarrow 0$, but it does exist for a finite, nonzero resolution $l$. After applying (2), the distance between two points in the Koch curve can be estimated and is given by
$d(l)=\int_{a}^{b}\,\sqrt{1+(\partial_{x}F(x,l)})^{2}\,dx\,,$ (3)
where $a$ and $b$ are the $x$-coordinates of the two points in the Koch curve, and $\partial_{x}$ denotes derivation with respect of $x$.
The length of a segment on the Koch curve depends on the smoothing function $h$ and the scale parameter $l$. We shall refer to this interplay between the resolution and non-differentiability as “fractality” in a broad sense. In Sec.3, we shall show that something similar can happen in QFT, where the role of $f$ and $F$ is played by the fields before and after smoothing, respectively, and the role of $d(l)$ is played by the action functional.
2.1 Introducing a scale in Euclidean Green equation: Dimensional Reduction
Let us now move to the case of a $\mathbb{R}^{d}$ non-smooth, Euclidean manifold. Similarly to the function $F$ in (2), the Green function has to depend on a scale that we call $M$, with the dimension of a mass-energy. The physical interpretation is that the space appears smooth at energy much below $M$, but it starts to show its fractal features at higher energy, comparable with $M$. This may be either an intrinsic scale of Nature or an effective energy emerging from a deeper theory than QFT, a theory that should include gravity at the quantum level  [^2] Fractal properties induced by quantum gravity effects are discussed in [20]..
Thus the standard Green equation,
$\left(\partial_{\mu}\partial^{\mu}-m^{2}\right)G(x-y)=-\delta^{(d)}(x-y)\,,$ (4)
becomes
$\left(\partial_{\mu}\partial^{\mu}-m^{2}\right)G(x-y,M)=-g(x-y,M)\,,$ (5)
where $g(k,M)$, represents the smoothing of the Dirac’s delta since the left-handed side of (5) now depends on $M$. The standard case is reproduced for $M\rightarrow\infty$. For simplicity, we choose $g$ in the Gaussian form,
$g(x,M):=\left(\frac{M^{2}}{4\pi}\right)^{\frac{d}{2}}\,e^{-\frac{1}{4}M^{2}x^{2}}\,.$ (6)
The Fourier transform of 5 reads,
$\left(p^{2}+m^{2}\right)G(p,M)=e^{-\frac{p^{2}}{M^{2}}}\,,$ (7)
leading to
$G(p,M)=\frac{e^{-\frac{p^{2}}{M^{2}}}}{p^{2}+m^{2}}\,.$ (8)
Next, let us do the inverse-Fourier transform of (8) in $O_{d}$ symmetry and in the massless limit,
$G(s,M)=2^{d-2}\pi^{d/2}s^{2-d}\left(\Gamma\left(\frac{d}{2}-1\right)-\Gamma\left(\frac{d}{2}-1,\frac{M^{2}s^{2}}{4}\right)\right)\,,$ (9)
being $s=\sum_{i}^{d}x_{i}^{2}$, $\Gamma(x)$ the gamma function, and $\Gamma(a,x)$ the incomplete gamma function.
For the topological dimension $d>2$ and $M\rightarrow\infty$, one knows that the standard result, $G(s)$, diverges as $s^{2-d}$ for $s\rightarrow 0$. In contrast, the result in (9) is finite as $s\rightarrow 0$. We want to encode this degree of divergence or convergence into an estimator of the effective dimension. To this aim, it is convenient to define $R:=s^{-\frac{2-d}{d}}$, such that $G(R)\sim R^{d}$, and thus the logarithm derivative of $G$ gives,
$\frac{d\log(G(R))}{d\log R}=d\,.$ (10)
This equation assigns ad hoc an effective dimension $d$, equivalent to the topological dimension, in the standard case. The reason is to define the effective dimension by replacing $G(R)\rightarrow G(R,M)$,
$d^{eff}:=\frac{d\log(G(R,M))}{d\log R}=d-\frac{2^{3-d}de^{-\frac{1}{4}M^{2}s^{2}}(Ms)^{d-2}}{(d-2)\left(\Gamma\left(\frac{d}{2}-1\right)-\Gamma\left(\frac{d}{2}-1,\frac{M^{2}s^{2}}{4}\right)\right)}\,,$ (11)
where in the last step we have rewritten the result in terms of the original variable, $s$.
The expression in (11) greatly simplifies in the phenomenological case for which the topological dimension is four, $d=4$,
$d^{eff}|_{d=4}=4-\frac{M^{2}s^{2}}{e^{\frac{M^{2}s^{2}}{4}}-1}\,.$ (12)
Therefore, $d^{eff}$ continuously flows from 4, at low energy, to $d^{eff}=2$ at $r\approx 2.2M^{-1}$, and it asymptotically reaches $d^{eff}=0$ at infinite energy. The interpretation is that $d^{eff}$ is an energy-dependent estimator of the non-constant dimension of the space. Specifically, the quantity $d^{eff}$ may be interpreted as an estimator of the “singularity spectrum” defining spacetime as a multi-fractal manifold [21, 22] – see [23, 24] for other applications to physical models. It is worth stressing that for energy close to $M$, the effective dimension becomes two, in agreement with many, if not all, suggestions from approaches to quantum gravity – see [25] for a comprehensive review. At energy much larger than $M$, the effective dimension vanishes. At energies much larger than the Planck energy, the effective dimension vanishes, which may not be immediately apparent. However, if we assume that $M$ is comparable to the Planck energy, we can speculate that the space-time dimensionality loses its meaning at energies where quantum fluctuations of space-time become dominant. This aligns with the analysis based on the spectral dimension estimator [26].
Before attempting the canonical quantization in this scenario, in the next section, let us summarize and comment on the picture we have introduced. First, the spacetime has topological dimension $d$ (e.g., $d=4$), thus one calculates in the standard way any algebra, vector, tensor, etc. usually appearing in QFT. Second, we conceptualize that quantum fields perceive an effective (varying) dimension $d^{eff}<d$, and this is defined by the behavior of the propagator for large momenta. The highlighted scenario shares some features with the one in [16], being one main difference that we consider and motivate a varying effective dimension. Another distinction is that we face the problem of rendering consistent canonical quantization in QFT, thereby dealing with well-known constraints.
3 Canonical quantization with varying effective dimension
We now turn our attention to actual QFT (in 1+3 topological dimensions). The aim is to generalize the canonical quantization of fields to the case of a fractal space-time with a running effective dimension.
Consider a real scalar field $\phi(x)$ and write the Feynman propagator,
$G_{F}(x-y)=\langle 0|\mathcal{T}\phi(x)\phi(y)|0\rangle\,,$ (13)
with $\mathcal{T}$ being the time-ordering operator.
In the same spirit as section 2.1, assume that space-time has fractal properties at scale $M$ and thus $G_{F}(x-y)$ has to be smoothed as $G_{F}(x-y,M)$. From (13), it follows that the field $\phi$ must depend on the parameter $M$. One can visualize the field $\phi(x,M)$ in terms of a convolution with a smoothing function, parametrized by $M$, as in (2).
As in the usual QFT, it is convenient to write the classical field, before quantization, in the Fourier representation (fixing $d=4$)
$\phi(x,M)=\frac{1}{(2\pi)^{4}}\int\phi(p,M)e^{ipx}d^{4}p\,.$ (14)
Essentially, the standard quantization consists of replacing $\phi(p)$ with the ladder operators, $\phi(p)\rightarrow a(\vec{p}),a^{\dagger}(\vec{p})$ (from here on, we denote the 3-vector with $\vec{x}$ to distinguish it from $x$, denoting the 4-vector). In our case, we first recast $\phi(p,M)$ as
$\phi(p,M)=\sqrt{r(p,M)}\hat{\phi}(p)\,,$ (15)
where $r$ is a positive definite function, and $r\approx 1$ for $p\ll M$, reproducing standard QFT at low energy, but changing physics at deep UV. Then we have  [^3] To keep contact with [16], this expression can be regarded as a Stieltjes-Fourier transform of measure $d\mu_{H}:=\sqrt{r(p,M)}d^{4}p$.,
$\phi(x,M)=\frac{1}{(2\pi)^{4}}\int\sqrt{r(p,M)}\hat{\phi}(p)e^{ipx}d^{4}p\,.$ (16)
Next, we promote $\hat{\phi}(p)$ as an operator, finally writing it in terms of the ladder operator. Specifically, we have,
$\phi(x,M)=\frac{1}{(2\pi)^{3}}\int\sqrt{\frac{r(p,M)}{2\omega_{p}}}\left[a(\vec{p})e^{-ipx}+a^{\dagger}(\vec{p})e^{ipx}\right]d^{3}\vec{p}\,,$ (17)
where the standard dispersion relation is implemented into $r(p,M)$:
$r(p,M):=r(p_{0}=\omega_{p},\vec{p},M)\,,$ (18)
with $\omega_{p}=\sqrt{\vec{p}^{2}+m^{2}}$. As it will be clear later, this is a necessary condition.
One implements the standard canonical commutation relations (CCR) but on the field $\hat{\phi}$ in (15),
$[a(\vec{p}),a^{\dagger}(\vec{p^{\prime}})]=\delta(\vec{p}-\vec{p^{\prime}})\hskip 10.00002pt\Leftrightarrow\hskip 10.00002pt[\hat{\phi}(t,\vec{x}),\hat{\pi}(t,\vec{y})]=i\delta(\vec{x}-\vec{y})\,,$ (19)
where $\hat{\pi}=\dot{\hat{\phi}}$ is the conjugate variable of $\hat{\phi}$, and the square brackets denote the commutator.
One may also define,
$a(\vec{p},M)=\sqrt{r(\vec{p},M)}a(\vec{p})\,,$ (20)
such that CCR can be rewritten as,
$[a(\vec{p},M),a^{\dagger}(\vec{p^{\prime}},M)]=r(\vec{p},M)\delta(\vec{p}-\vec{p^{\prime}})=r(p,M)\delta(\vec{p}-\vec{p^{\prime}})\,,$ (21)
where the last equality comes from (18). The action on the vacuum becomes
[…]
Therefore, the action of the field on the vacuum leads to a plane wave but with a scale-dependent prefactor:
$\phi(x,M)|0\rangle=\frac{1}{(2\pi)^{3}}\int\frac{d^{3}p}{\sqrt{2\omega_{p}}}\sqrt{r(\vec{p},M)}e^{ipx}|1_{\vec{p}}\rangle\,.$ (23)
The generalization of the quantization, in the presence of dimensional reduction, of vector and fermion fields is straightforward. It is sufficient to decompose the fields as in (15) and then write covariant CCR, for vector fields, and anti-commutation relations (ACR) for fermions. The Reader might worry about the quantization of vector fields, then of gauge bosons, in the presence of a dimensional scale ($M$) since this may resemble a cutoff, thus inconsistent with gauge invariance. However, this is not the case. A sharp cutoff rules out the high momentum modes breaking the gauge invariance, conversely, the Fourier transforms of the fields, e.g., (17), ranges up to infinite energy.
3.1 Propagator
Generalizing (13), the Feynman propagator is
$G_{F}(x-y,M)=\langle 0|\mathcal{T}\phi(x,M)\phi(y,M)|0\rangle\,.$ (24)
Replacing the field with the representation in (17) and using (23) gives,
$G_{F}(x-y,M)=\frac{1}{(2\pi)^{3}}\int d^{3}\vec{p}\frac{r(\vec{p},M)}{2\omega_{p}}e^{-ip(x-y)}\,,$ (25)
which has to be written in terms of 4-vectors. We achieve this in the usual way but with some attention to the function $r$. We call for shortness $s:=x-y$ and consider the integral,
[…]
Writing $p^{2}-m^{2}=(p-\omega_{p})(p+\omega_{p})$, integration on the known Feynman contour, and picking up the pole at $\omega_{p}-i\epsilon$ leads to,
$\frac{1}{(2\pi)^{3}}\int d^{3}\vec{p}\frac{r(p_{0}=\omega_{p},\vec{p},M)}{2\omega_{p}}e^{-ip(x-y)}\,.$ (27)
Comparing this with (25) implies,
$r(\vec{p},M)=r(p_{0}=\omega_{p},\vec{p},M)\,,$ (28)
which is $r(p,M)$ due to (18). Therefore, the propagator in momentum space, expressed through 4-vectors, is
$G_{F}(p,M)=i\frac{r(p,M)}{p^{2}-m^{2}+i\epsilon}\,.$ (29)
It is important to elaborate on the implications of (18) and (28). By comparing (8) and (29) (regardless that the latter is in Minkowski space) , one sees that $r$ is an even function in momentum. This will turn out to be problematic for the consistency of the theory. The function $r$ must be odd, so we can modify it as
$r(p,M,v)=V\,e^{-(\frac{p}{M}-v)^{2}}$ (30)
where $v$ is a constant 4-vector necessary to build $r$ as a function not only of $p^{2}$, and $V=\exp[v^{2}]$ is a normalization constant.
Among others, the introduction of a special direction, $v$, will imply the breaking of the rotational invariance of the vacuum. Even more important for the rest of this work, the form of the function $r$ in (30), odd in the variable $p$, will break the spatial translational invariance of the vacuum, with relevant impact for the consistency of the theory, discussed in section 4.
However, the insight on dimensional reduction from (8) remains unchanged since it is only based on arguments of convergence or divergence of the propagator, dominated by the $p^{2}$ contribution, which reflects on the loop behavior or the couplings running. Notice also that $r$, in the form of (30), does not trivialize on-shell yielding effects even for tree-level processes.
3.2 Translations, rotations, and the vacuum
Let us first focus on the spatial translations which, as anticipated, play a central role in the proposed framework.
Similarly to the standard case, the 3-momentum operator is given by  [^4] The vacuum-to-vacuum expectation value of $P_{i}$ is zero in the standard QFT, but a priori not the one of $P_{0}$. Often, this is artificially made zero by Normal Product. One reason is that the corresponding integral rapidly diverges. This is in contrast with the QFT in the multifractal spacetime, where the finiteness does not require artifacts in the regularization.
$P_{i}=\int d^{3}\vec{x}\,\dot{\phi}(x,M)\partial_{i}\phi(x,M)=\frac{1}{(2\pi)^{3}}\int d^{3}\vec{p}\,\frac{p_{i}}{2}\left(a^{\dagger}(\bar{p},M)a(\bar{p},M)+a(\bar{p},M)a^{\dagger}(\bar{p},M)\right)\,,$ (31)
being the dot the temporal derivative.
By evaluating the vacuum expectation value with the help of equation (22), one finds (for shortness, we indicate $r(p,M,v)$ just as $r(p)$)
$\langle 0|P_{i}|0\rangle=\frac{1}{(2\pi)^{3}}\int d^{3}\vec{p}\,\frac{p_{i}}{2}r(\vec{p})\,,$ (32)
which generalizes the standard expression, here modified by $r$.
Due to the form of the function $r$ in equation (30), and in contrast with the standard QFT, this integral is non-zero but finite. In other words,
$P_{i}|0\rangle\neq 0\,,$ (33)
namely, the vacuum possesses momentum. The $P_{i}$ are the generators of the spatial translation, $T=\exp[i\,b_{j}P_{j}]$, being $b_{j}$ the translational parameters, so the vacuum is not translational invariant:
$T|0\rangle\neq|0\rangle\,.$ (34)
Since rotations are written as volume integrals of the type in equation (31), where one takes the textbook expressions and replaces $\phi(x)\to\phi(x,M)$ and then $a(\vec{p})\to a(\vec{p},M)$, the analog conclusion of equation (33) also holds for rotations. Thus the theory, due to the deformation of the generators of translation or rotations caused by $r$, shows a breaking at the quantum level of the Poincaré invariance.
4 Consistency of the theory
In this section, our proposal confronts two milestones of QFT: the Kallen-Lehmann representation of the propagator; the Gell-Mann and Low theorem for the propagator.
While the nonperturbative Kallen-Lehmann representation bounds the behavior of the propagator, thus providing a consistency check for the theory, the Gell-Mann and Low formula, base for perturbation theory, is not a test for the theory but is an improved result within nonstandard QFT including the scale $M$.
Notice that both these well-known results are related to the Poncaré (non)invariance of the vacuum. We shall now go through the issue in detail.
Kallen-Lehmann spectral representation.
In standard QFT, the Kallen-Lehmann representation of the propagator reads,
$G(p)=\int_{0}^{\infty}d(\mu^{2})\rho(\mu^{2})\frac{1}{p^{2}-\mu^{2}}\,,$ (35)
where $\rho$ is a positive definite function, known as spectral function.
The above equation implies that the propagator cannot decay faster than $p^{-2}$ for large momentum, in apparent contrast with (29). However, the Kallen-Lehmann representation does not exist in the fractal QFT and this invalidates the above constraint. The existence of the Kallen-Lehmann representation in (35) requires the Poncaré invariance of the vacuum to manipulate the expression $\langle\Omega|\phi(x)\phi(y)|\Omega\rangle$. In particular, one has in the fractal QFT,
$\langle\Omega|\phi(x)\phi(y)|\Omega\rangle=\sum_{n}\langle\Omega|\phi(x)|n\rangle\langle n|\phi(y)|\Omega\rangle=e^{ip_{\Omega}(x-y)}\sum_{n}e^{-ip_{n}(x-y)}|\langle\Omega|\phi(0)|n\rangle|^{2}\,,$ (36)
In the first step, we used the completeness relation for a complete set $|n\rangle$, eigenvalues of the four-momentum, in the last step, we implemented the non-translational invariance of the vacuum, via the vacuum momentum $p_{\Omega}$. However, this is infinite since one has a finite momentum density in (32) and (33), corresponding to an infinite momentum for an infinite volume. Therefore, the expression in (36) becomes ill-defined.
We thus conclude that the Kallen-Lehmann representation of the two-point function does not exist if the vacuum is not invariant under translations, and therefore it does not constrain the expression (29) to behave as $1/p^{2}$ for large $p$.
Gell-Mann and Low theorem and S-matrix construction.
While the previous paragraph is a consistency check, this section demonstrates how fractal QFT solves some fundamental issues in standard QFT through Poincaré non-invariance of the vacuum. This property, shown in equation (33), enables the construction of a consistent $S$-matrix and proper application of the Gell-Mann and Low formula since broken translational invariance invalidates Haag’s theorem [17]. Standard QFT suffers from Haag’s theorem [17], which prohibits a non-trivial interaction picture or perturbation theory – a no-go result often overlooked in contemporary literature.
Let us formulate this precisely. We denote the vacuum states in interactive and non-interactive Fock spaces as $|\Omega\rangle$ and $|0\rangle$, respectively, and represent interactive and non-interactive fields as $\phi$ and $\phi_{0}$. The evolution operator relating free and interacting theory for $t>t^{\prime}$ is:
[…]
Standard QFT assumes the existence of the unitary operator connecting the free and interactive fields through:
$\phi(t,\vec{x})=U^{\dagger}(t,t_{0})\phi_{0}(t,\vec{x})U(t,t_{0})\,.$ (38)
Moreover, the free and interactive vacuua are related as (calling $t^{\prime}=0$ and $t=T$, as $T\rightarrow\infty$) [27],
$|\Omega\rangle=\frac{U(0,\pm T)|0\rangle}{\langle 0|U(0,\pm T)|0\rangle}\,.$ (39)
Plugging (39) into $\langle\Omega|\mathcal{T}\phi(x)\phi(y)|\Omega\rangle$ and performing simple algebra yields the Gell-Mann and Low relation:
[…]
Unfortunately, Haag’s theorem establishes that equations (38) and (39) are valid only in the free-field case, namely, when $U$ is the identity – meaning that the interaction picture exists only when no interaction exists  [^5] For an overview of the interpretations of Haag’s theorem, see [28]..
For the present discussion, the crucial point is that fractal QFT circumvents this issue through its non-translational invariant vacuum. The proof of Haag’s theorem depends on vacuum translational invariance (see appendix A), a condition explicitly violated in fractal QFT as shown in equation (33). This violation enables fractal QFT to construct a consistent operator $U$ in (38) and (39), while the logical steps to arrive at (40) remain unchanged. As a result, fractal QFT guarantees a robust perturbative framework, remarkably improving the standard theory – see also section 5.
5 Possible Implications and Discussion
In this section, we explore some implications of fractal QFT. The most striking consequence is perturbative finiteness – all loops become convergent. Moreover, we argue that perturbative series behave better at large order $n$ due to the absence of UV renormalons. Recall that renormalons lack any semi-classical interpretation that drives the factorial ($n!$) divergence in renormalized perturbation theory [29].
Loop finiteness.
The appearance of the function $r$ in (29) renders loop integrals finite, creating a perturbatively finite theory. Consider the 4-point correlator in the $\phi^{4}$ model with interaction Lagrangian $\frac{\lambda}{4!}\phi^{4}$—the so-called ”fish-diagram.” This becomes finite due to the suppression from $r$ at high momentum. Since $r(p)\approx 1$ for $p\ll M$ while providing exponential suppression for $p>M$, the integral behaves as $\log(p^{2}/\mu_{0}^{2})$ for $p\ll M$, and as $M^{2}/p^{2}\exp(-p^{2}/M^{2})$. Consequently, the one-loop running becomes:
$\left\{\begin{array}[]{ll}\lambda(\mu)\simeq\lambda(\mu_{0})+\left(1+\lambda(\mu_{0})\beta_{1}\log(\mu^{2}/\mu_{0}^{2})\right)&\mu<M\\[10.0pt] \lambda(\mu)\simeq\lambda(M)\left(1-\lambda(M)\beta_{1}\frac{M^{4}}{4\mu^{4}}e^{-\frac{\mu^{4}}{M^{4}}}\right)&\mu\geq M\,,\\ \end{array}\right.$ (41)
where the usual one-loop factor is $\beta_{1}=3/(16\pi^{2})$. The first line of (41) shows standard running; the second reveals that the coupling quickly reaches an asymptotic constant value for $\mu\geq M$.
In summary, any loop constructed in the model becomes finite, and no Landau pole exists due to (41).
We should also examine modifications to on-shell processes, such as $2\rightarrow 2$ scattering. At leading order in $\lambda$ expansion, we have:
$\langle f|S^{(1)}|i\rangle=\langle 1_{p},1_{p^{\prime}}|(-i\,\lambda)\,\phi^{-}(x)\phi^{-}(x)\phi^{+}(x)\phi^{+}(x)|1_{k},1_{k^{\prime}}\rangle=$
$\frac{1}{(2\pi)^{8}}(-i\,\lambda)\,\delta^{(4)}(p+p^{\prime}-k-k^{\prime})\,\frac{\sqrt{r(p)\,r(p^{\prime})\,r(k)\,r(k^{\prime})}}{\sqrt{(2\omega_{\vec{p}})\,(2\omega_{\vec{p}^{\prime}})\,(2\omega_{\vec{k}})\,(2\omega_{\vec{k}^{\prime}})}}\,.$ (42)
The superscripts -- and $+$ denote the creation and destruction operator parts of the field. The standard result is modified only by the factors $r$ in (42). Given the form of the function $r$ in (30), the part in $p^{2}$ becomes trivial onshell, $p^{2}=m^{2}$, while the part in $p_{\mu}v^{\mu}$ produces an anisotropy. Some momentum directions may be enhanced, depending on the constant four-vector $v$. Of course, all these modifications are suppressed at energies much lower than $M$.
Absence of the UV renormalons.
We have discussed the perturbative finiteness of the theory above, now let us consider what to expect at the non-perturbative level. Renormalons [29, 30] can be interpreted as a bridge between perturbative and non-perturbative physics. The singularities in the Borel transform due to renormalons lie on the semi-positive axis and hamper any Borel-Laplace resummation. They are not related to any semi-classical expansion and are considered genuine non-perturbative objects related to renormalization [30]. In the literature, their presence represents a failure of the perturbative renormalization procedure since the resulting series cannot be resummed without ambiguities [29].
Returning to the $\phi(x)^{4}$ model, renormalons relate to factorially divergent series [^6] Independently of the instantons that notwithstanding leading to a $n!$ large order behavior of the perturbation theory series, their existence does not imply any inconsistency in the theory, as explained in [29]. obtained by evaluating Feynman diagrams with many insertions of the same sub-diagram (the fish-diagram). The origin of the factorial $n!$ (where $n$ is the perturbation theory order) stems from the logarithmic high-energy behavior of the fish-diagram (sub-diagram) inserted $n$ times. While we refer to [31] for a complete derivation of the $n!$ renormalon behavior (see [32] for an alternative approach), here we illustrate its origin with a simple example. Consider the integral:
$\int_{\mu_{0}}^{\mu}\log(\frac{p}{\mu_{0}})^{n}d^{4}p\,.$ (43)
This expression mimics a so-called bubble diagram, i.e. a loop integral with $n$ nested occurrences of the 4-point correlator (fish diagram), which has logarithmic UV behavior in standard QFT.
Renormalons are calculated from the finite part of the loop, so focusing on the finite part of (43) as $\mu\rightarrow\infty$, we obtain:
$\left(-\frac{1}{4}\right)^{1+n}\mu_{0}^{4}\,n!\,.$ (44)
This suffices to trace the factorial divergence of the series due to renormalons back to the logarithmic UV behavior of the running coupling.
Here lies the improvement provided by dimensional reduction theory. Due to (41), the behavior for $\mu\rightarrow\infty$ is not logarithmic, unlike standard QFT. This suffices to eliminate the renormalon ambiguities.
It is important to emphasize that the absence of UV renormalons and the absence of the UV Landau pole are distinct phenomena that need not occur together. Typically, one might encounter situations where renormalons emerge from the perturbative evaluation of bubble diagrams even when the Landau pole is absent due to nonperturbative dynamics, as discussed in [33, 34]. However, our current framework provides a more comprehensive solution: within the perturbative approach itself, both the UV Landau pole and renormalons simultaneously disappear, highlighting the theoretical consistency of our model.
A comment is finally in order. In [35], we conjecture that since UV renormalons indicate a failure of self-consistent perturbative renormalization, they might be reinterpreted as ignorance of the no-go provided by the Haag theorem. Standard perturbative divergences might also signal a warning about this no-go, though perturbative renormalization addresses this problem. However, renormalized perturbation theory needs resummation for self-consistency, which renormalon ambiguities prevent. The present work supports this interpretation: we show that when the Haag theorem’s no-go disappears, so do renormalons (and perturbative divergences).
Implications for the standard model.
Within the multifractal spacetime, the standard model becomes automatically finite at any order in perturbation theory and avoids UV Landau poles. Consequently, it achieves asymptotic safety and remains valid at all scales—see [36] for an in-depth discussion on the asymptotic safety paradigm in QFT. For example, figure 1 illustrates the one-loop running of the standard model’s gauge couplings.
It’s worth commenting on the nature of asymptotic safety shown in Fig. 1. Wilson first argued for the necessity of a non-trivial UV fixed point for theoretical consistency [37]. In our case – QFT with diminishing dimension – the UV fixed point is reached only at infinity. Nevertheless, the couplings remain finite and possibly small. While the fixed point is attained at infinite energy, the running exponentially converges to a constant value. Additionally, as explained previously for the $\phi^{4}$ model, the absence of UV renormalons enhances perturbation theory’s robustness.
Last but not least, as illustrated in (42), the multifractal QFT predicts direction-dependent cross-sections. Phenomenologically, within the standard model, this would probably be the most striking “smoking gun” signal of the underlying space-time fractality. In parallel, the Lorentz violating aspects of the proposed scenario should be phenomenologically bounded, being this kind of studies an active research field [38]. A detailed phenomenological analysis is beyond the scope of this work.
Figure 1: Running of the gauge couplings of the standard model, which asymptotically reach constant values.
6 Conclusions and outlook
Motivated by the possibility that spacetime’s effective dimension may evolve with energy scale, we have explored this concept from a quantum field theory perspective, particularly within canonical quantization. Crucially, consistency of the quantum field theory points to a quantum nature of spacetime, with classical smoothness emerging only as a low-energy approximation.
We have shown how to quantize fields in a spacetime with multifractal properties and how this automatically implies dimensional reduction, ensuring consistency with established field theory constraints. Notably, not only is quantization compatible with a varying dimension, but such variation significantly enhances QFT’s robustness, enabling rigorous $S$-matrix construction in the interaction picture (perturbation theory). This results in a finite theory without loop divergences and improves the perturbative series’ behavior. The significance of our contribution lies in demonstrating that non-differentiable spacetime structures are not only compatible with quantum field theory but actively improve its mathematical foundation. All this is achieved while retaining all known behaviors and predictions of standard QFT at low energies, yet predicting new behaviors at large energies, mainly through “asymptotic safety” and small anisotropies at high energy scatterings.
When comparing our work with existing literature, some apparent similarities arise between our quantization in section 3 and the works in [39, 40], where the author proposes scale-dependent QFT. However, the concept of scale dependence in those works has no connection to our fractality and dimensional reduction. The conceptual framework, technical details, and outcomes are fundamentally different. Some analogy in intent exists with our earlier proposal [41]. However, important differences exist concerning what we proposed in this work. The approach in [41] requires modifications due to defining the action via Stieltjes integrals, entailing modifications to the equations of motion that must be addressed with certain constraints. The Stieltjes-integral action approach in [41] is more technically involved and becomes inconsistent when introducing Lagrangians with additional fields, particularly fermions. Conversely, our present approach requires only a specific quantization of the standard action, incorporating a dimensional reduction scale, and it remains unaltered with the addition of new fields, fermions included. A radically different route to model fractal spaces is the one in [42], based on fractional calculus. A final parallel can be drawn with [43, 44]; however, these references attempt to address varying topological dimensions, rendering the framework genuinely distinct.
Looking beyond particle physics, we should also comment on the possible impact of (30) on the standard cosmological model. First, the vacuum energy density becomes finite, like any integral in the theory, but remains remarkably large – approximately $O(M^{4})$. Standard QFT requires regularization and renormalization to control divergences; conversely, in fractal space-time QFT, integrals are inherently finite, though the renormalization group equation remains applicable. This property is generally independent of divergences. Thus, even in a universe with running dimensions, QFT does not produce a small cosmological constant. The smallness must instead be understood within the renormalization group framework, where parameter values are in principle arbitrary and cannot be predicted.
Second, and more intriguingly, the introduction of a specific direction – represented by vector $v$ in (30) – indicates a fundamental anisotropy of space-time. As highlighted in (33), in QFT with dimensional reduction, the vacuum possesses not only energy but also momentum along a particular direction. Consequently, even though all the possible new effects are suppressed by the scale $M$, the stress-energy tensor is expected to develop non-diagonal components. This might impact the Friedmann equation, potentially suggesting non-standard cosmological effects, namely, departing from the cosmological principle [45]. Further exploration of this direction is compelling but beyond our current scope.
In summary, while phenomenological studies in the literature investigate potential effects of varying dimensions [46], we believe our work provides the theoretical foundation for such scenarios. A dedicated study will be needed to analyze further consequences for particle physics and cosmology.
Acknowledgement
AM thanks Emanuela Pichelli for discussions during this work on the appearance of fractals in the physical world.
Appendix A S-matrix and Poincaré invariance of the vacuum
This appendix highlights some known problems in building a consistent $S-$matrix in interaction picture in QFT.
Equation (40) relies on the interaction picture where the free field, $\phi_{0}$, and the interactive field, $\phi$, are unitarly related:
$\phi(t,\vec{x})=U(t_{0},t)^{\dagger}\phi_{0}(t,\vec{x})U(t_{0},t)\,,$ (45)
and the vacua $|0\rangle$ and $|\Omega\rangle$ of the free and interactive Fock spaces, respectively, are two distinct vectors, e.g., as (39),
$|\Omega\rangle\neq|0\rangle\,.$ (46)
In standard QFT, namely, in the absence of the fractality and reduction scale, the unitary operator $U(t_{0},t)$ exists only in the trivial case, with no interaction, for which such an operator is just the identity. This result is known as the Haag theorem, heavily relying on the translational invariance of the vacuum. It may be helpful to recall how this invariance leads to the no-go of the theorem, following [35] – see [47] for a rigorous proof within axiomatic QFT.
For shortness, we denote $U(t_{0},t):=U$ and call $T_{0}$ and $T$ the translational operator in the free and interactive Fock spaces. The translational invariance of the vacuua reads,
$T_{0}|0\rangle=|0\rangle\hskip 20.00003ptT|\Omega\rangle=|\Omega\rangle\,.$ (47)
Calling $\vec{b}$ the translational parameter, the operators $T_{0}$ and $T$ act on the fields as,
[…]
From the latter and (45), we have,
$T^{\dagger}\phi(t,\vec{x})T=\phi(t,\vec{x}-\vec{b})=U^{\dagger}\phi_{0}(t,\vec{x}-\vec{b})U=U^{\dagger}T_{0}^{\dagger}\phi_{0}(t,\vec{x})T_{0}U\,.$ (49)
On the other hand, we have,
$T^{\dagger}\phi(t,\vec{x})T=T^{\dagger}U^{\dagger}\phi_{0}(t,\vec{x})UT\,.$ (50)
Comparing (49) and (50) yields,
$UT=T_{0}U\,.$ (51)
Next, multiply from the right (51) for $|\Omega\rangle$,
$UT|\Omega\rangle=U|\Omega\rangle=T_{0}U|\Omega\rangle\,.$ (52)
and comparing the last equality with the first of (47) (i.e. $T_{0}|0\rangle=|0\rangle$) gives
$U|\Omega\rangle=|0\rangle\,,$ (53)
or equivalently
$|\Omega\rangle=U^{\dagger}|0\rangle\,.$ (54)
These equations conflict with (39).
Moreover, from (53) and (54), we can write,
$\langle\Omega|U|\Omega\rangle=\langle\Omega|0\rangle\hskip 20.00003pt\langle 0|U^{\dagger}|0\rangle=\langle 0|\Omega\rangle\,.$ (55)
Finally, these two equations lead to the following chain of equalities,
$\langle\Omega|U|\Omega\rangle=\langle\Omega|0\rangle=\left(\langle 0|\Omega\rangle\right)^{\dagger}=\langle 0|U|0\rangle\,,$ (56)
implying (modulo an irrelevant overall phase)
$|\Omega\rangle=|0\rangle\,,$ (57)
in contradiction with (46), unless the unitary operator $U$ coincides with the identity. This case, however, is the one with no interactions, hence the Haag theorem. Since the $S-$matrix is $U(-\infty,\infty)$, the Haag theorem is a no-go for perturbative QFT. The multifractality and dimensional reduction dramatically change (47), enabling the construction of a consistent $S-$matrix.